EE 211 Quiz 1  Fall 1998

EE 211 Section 1 Course Overview

1. KVL, KCL, Element Equations

(a) For a circuit with N elements, get 2N equations with 2N

unkowns (i_{k} and v_{k}).

2. Element Equations

(a) Independent Sources i. Voltage Source ii. Current Source

(b) Resistor

(c) Open Circuit

(d) Short Circuit

(e) Switch

(f) Dependent Sources i. VCVS ii. VCCS iii. CCVS iv. CCCS

(g) Capacitor

(h) Inductor

EE 211 Quiz 1  Fall 1998

3. Techniques

(a) Equivalent Circuits i. Resistors in parallel and series ii. Capacitors in parallel and series iii. Inductors in parallel and series iv. Source Transformations v. Thev́enin Equivalent vi. Norton Equivalent

(b) Circuit Reduction

(c) Voltage and Current Division

(d) Superposition

(e) Unit Output

(f) Node-voltage

(g) Mesh-current

4. Power Transfer

5. Active Circuit Analysis

6. Op-Amp Analysis (a) Node-voltage

(b) Building Blocks

(c) Block Diagrams

7. Input and Output Resistance

EE 211 Quiz 1  Fall 1998

8. Signals (a) Impulse

(b) Step

(c) Ramp

(d) Exponential

(e) Sinusoidal

(f) Combined i. Sum of two exponentials ii. Damped exponential iii. Exponential times ramp

9. DC Analysis with capacitors and resistors

10. Differential equations (a) Characteristic equation

(b) Natural (transient) and forced (steady-state) responses

(c) Zero-input and zero-state responses

11. First order circuits (a) y(t) = (IC − F C)e^{−t/T C} + F C

12. Second order circuits (a) Underdamped, overdamped, critically damped